OpenStudy (anonymous):

Darla is building a new desk. To make sure she had made a square corner she measures 7ft from the corner along one edge and 5ft along the other edge. How long should the diagonal be between those two pants if the corner is a right angle?

OpenStudy (anonymous):

Pythagoream Theorem A^2 + B^2 = C^2 25 + 49 = C^2 Diagonal Should Be sqrt{74}, or about 8.62 Units